Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $180 of her medical bills, plus 30% of the rest. Under plan B, Giselle would pay the first $200, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $200 in bills.
Giselle would save with plan B if she had more than $540 in bills. Is this correct?

Answer by stanbon(75887)

Bayside Insurance offers two health plans.
Under plan A, Giselle would have to pay the first $180 of her medical bills, plus 30% of the rest.
---
A(x) = 180+0.30(x-180)
------
Under plan B, Giselle would pay the first $200, but only 20% of the rest.
---
B(x) = 200+0.20(x-200)
---
For what amount of medical bills will plan B save Giselle money? Assume she has over $200 in bills.
----
Solve B(x) < A(x)
200+0.2x-40 = 180+0.3x-54
160 + 0.2x = 126+0.3x
0.1x = 34
x = $340
Plan B saves money if bills are greater than $340
===================================================
Cheers,
Stan H.
===============